# Chlorophyll-a Mapping and Prediction in the Mar Menor Lagoon Using C2RCC-Processed Sentinel 2 Imagery

### v1

@@ II-B Related Work
where $R(\lambda_{i})$ corresponds to the different band reflectances. During the study period from May 2015 to May 2017, the best results were achieved using the index normalized with the green and blue bands, reaching an R² of 0.88. Further work from the same authors [25] focused on obtaining maps using the empirical algorithm previously proposed, and they also studied the relationship between Chl-a and turbidity. The study revealed the intricate variability between these parameters, as they were not consistently correlated due to factors such as atmospheric conditions, water column movements, and phytoplankton distribution.
@@
In 2020, [26] used machine learning (ML) models to predict the concentration of Chl-a in the Mar Menor lagoon based on physical and chemical water data. Two approaches were evaluated: multilayer neural networks and support vector regression (SVR). Parameters such as water temperature, pH, suspended solids, and turbidity were used as predictor variables. However, this study does not use remote sensing or spectral bands; it is based exclusively on in situ measured parameters, so there are no spectral formulas or satellite image use. Yet, the study emphasizes that ML models are especially useful for monitoring this type of environment.
@@
The literature regarding remote sensing with Sentinel 2 for Mar Menor starts to proliferate in 2021. [27] proposed a methodology using ML and deep learning (DL) models to estimate surface (less than 0.5 m) Chl-a. The inputs included band combinations such as the normalized difference chlorophyll index (NDCI), the normalized difference turbidity index (NDTI), and $B3/B8$ ratios, among others. Images are resampled up to 60 meters, and the most important features in the optimal model are the $B3/B8$ ratio, NDTI, and bands 9 and 2. Additionally, feature selection was explored in three scenarios: all variables, removal of features with correlations greater than 0.75, and Principal Component Analysis (PCA). Overall, maintaining all variables yielded the best results.
@@
[28] combined the two previously mentioned satellites: Sentinel 2 and Landsat 8. Their study focused on a brief period of time in 2021 and they implemented atmospheric corrections with ACOLITE, a toolbox designed to correct Level 1 products over marine, inland and coastal waters. Then, they estimated Chl-a using the bio-optical algorithm OC3 [29] based on water inherent properties and applied the model to the whole lagoon. In the case of [30], rather than applying the model to each pixel to map the entire water surface, the best performing model was applied only to points where in situ measurements were taken. Then, an interpolation method such as distance weighting or kriging was used to map the rest of the water surface.
@@
In [31], SNAP was used to correct Sentinel 2 images with the set C2X of C2RCC. With the atmospherically corrected images, reflectances are grouped within a 3 x 3 pixel window. Then, Chl-a is estimated using empirical models in which different band combinations are tested to fit the line $y=ax+b$. The band combination that yielded the best results is:
@@
While previous studies focused only on surface chlorophyll (less than one meter), [32] explored prediction across the entire water column. They used several ML and DL models along with Sentinel-3 imagery to predict Chl-a concentrations at various depths for each in situ measurement point. They used 21 OLCI spectral bands as input and feature selection based on statistical significance. The best results are achieved on the surface, and as the depth increases, metrics deteriorate.
The most recent study in the Mar Menor region [33] focused on using MODIS, VIIRS, and Sentinel-3 sensors to monitor Chl-a. The study developed an empirical algorithm named BELA based on the red-to-green ratio, which is expressed by the following formula: $Chl=0.353RG^{3}+2.132RG^{2}+3.905RG+2.110$ where $RG=\log_{10}(R_{670}/R_{550})$. However, when we applied this relationship to Sentinel 2 reflectances and our data, the results were not satisfactory. This may be due to the significant differences between the sensors in this article and the MSI, as well as the atmospheric corrections we used.
Remote sensing has been also widely used in other areas with turbid and productive water bodies. [34] introduced the NDCI to estimate Chl-a in Mobile Bay and the Mississippi Delta with MERIS imagery adjusting the model with simulated data to cover a wide range of chlorophyll values. Building on this, [35] also employed NDCI in combination with probabilistic-based networks, arguing that chlorophyll is only one of several components influencing reflectance. [36] uses MODIS and proposes the idea of using a polynomial with logarithms for low values of Chl-a and a SVM for Chl-a greater than 3 mg/m3 while [37] explores the use several satellites in a tropical estuarine lagoon in Brasil. Finally, focusing on the Ebro Delta in Spain, [38] applied the C2RCC and C2X processors for atmospheric correction of Sentinel 2 imagery and tested several band combinations using linear, simple polynomial, and logarithmic models for chlorophyll estimation.
Table II summarizes all the articles discussed in a compressed format, so the reader can easily compared the methods used, band combinations, chlorophyll ranges, etc.
TABLE II: Related Work summary table. Units for RMSE and MAE are mg/m3
@@ 10 B1 to B12
RF, LBM, XGB - - No
[31] Mar Menor Sentinel 2 10
@@
[32] Mar Menor Sentinel 3 300 21 OLCI Bands
@@
[38] Ebro Delta Sentinel 2 10
@@
While numerous studies have demonstrated the potential of remote sensing for monitoring Chl-a in turbid and optically complex waters, certain aspects remain underexplored, particularly in regions such as the Mar Menor. Open-source methodologies that integrate long-term datasets, cover a broad range of Chl-a concentrations, and include systematic comparisons of multiple modeling strategies are still relatively scarce. Moreover, although neural network-based tools such as C2RCC and C2X are increasingly used, their integration into customized regional workflows and their comparison with other modeling approaches are areas that merit further development.
In this study, we address these gaps by developing and openly releasing a methodology tailored for the retrieval of Chl-a in optically complex coastal waters, using nearly a decade of Sentinel 2 data at 10 m resolution. The approach systematically evaluates a wide range of band combinations and modeling techniques, including ensemble learning and neural networks (C2RCC, C2X, and C2X-Complex) for atmospheric corrections, within a robust cross-validation framework. Additionally, we provide clear guidance and visual examples–in Section V and in the Github repository [^1] https://github.com/Antonio-MI/mar-menor-chl–of how the best-performing models can be operationally applied to generate Chl-a maps, contributing to improved transparency, reproducibility, and practical utility in regional water quality monitoring.
III Materials and methods
@@
III-A Study area
The Mar Menor is a coastal lagoon separated from the Mediterranean Sea by La Manga, with an approximate area of 135 km2. The region’s climate is classified as a dry Mediterranean climate, with an average annual temperature ranging from 17°C to 18°C. The summers are hot, with temperatures over 40ºC, and the winters are mild, with extremes in temperature mitigated by the influence of the nearby sea. The region experiences low rainfall, typically ranging from 270 millimeters to 350 millimeters per year [39]. The temperature of the water body is subject to variation according to the season. During the winter months, the temperature is approximately 15°C. In the spring, the temperature typically ranges from 16–18°C. During the summer months, the temperature reaches an average of 24-25°C, and there have been recorded temperatures of 31-32°C during recent episodes.
The salinity of the Mar Menor, classified as a hypersaline coastal lagoon, typically varies between 40 and 47 Practical Salinity Unit (PSU), which is considered high for a coastal lagoon and attributed to minimal rainfall and substantial evaporation. In hydrodynamic terms, the Mar Menor displays marked homogeneity in its internal behavior, with minor variations in level between sensors except during episodes of intense wind, where seiche-type oscillations are generated in the lagoon. The wind also generates a water level difference of up to 10 centimeters and directs surface currents in its direction, thereby facilitating water mixing and renewal [3].
III-B Methods
This subsection outlines the ML and DL models that were implemented. This selection explores a wide range of algorithms, covering classical statistical approaches and modern ML techniques. Some of these models were selected because previous studies have already demonstrated their suitability for predicting Chl-a and related water quality parameters in coastal and inland waters. Others were included to expand upon those findings and assess whether additional modeling paradigms could improve robustness and generalization. This includes distance-based methods, linear and regularized regression models, tree-based learners, and gradient boosting frameworks. Additionally, DL is represented, and ensemble strategies are used to combine individual predictors. This variety of models enables a thorough comparison of different learning paradigms and demonstrates their respective strengths in Chl-a prediction.
@@ III-B Methods
A supervised learning method in which the output is modeled as a linear combination of input features, capturing the approximately linear relationships between Chl-a concentration and spectral reflectance [40]. The model assumes a constant rate of change in the output for each unit change in the inputs, and its parameters are typically estimated by minimizing the sum of squared errors [41]. Implemented with LinearRegression from sklearn.
@@
A non-parametric regression method where the prediction for a given instance is based on the average of the target values of its $k$ closest neighbors in the feature space. In this work the weights are assigned proportionally to the inverse distance from the query point, and distance is computed as euclidean distance. KNN could suit this problem because it can capture local nonlinearities between reflectance and Chl-a concentration without assuming a specific functional form. Implemented with KNeighborsRegressor from sklearn [42].
@@
A regression algorithm derived from Support Vector Machines (SVM) that seeks to fit a function within an error margin (an epsilon-insensitive tube) while penalizing points that fall outside this margin. SVR effectively captures complex nonlinear relationships between environmental factors and Chl-a concentration, avoids overfitting through structural risk minimization, performs well with small sample sizes, and provides high prediction accuracy even when data are noisy [43]. Implemented using the SVR module from scikit-learn [44].
@@
A linear regression model that combines L1 (lasso) and L2 (ridge) regularization penalties. It is useful when there are many correlated features: L1 helps with feature selection (sparse coefficients), L2 stabilizes the coefficient values. In this context, ELN helps reduce redundancy between highly correlated spectral bands while maintaining model stability. Implemented with ElasticNet from sklearn [45].
Random Forest (RF) An ensemble learning method that builds many decision trees on bootstrapped subsets of the data and averages their predictions to reduce overfitting and improve generalization. RF can handle nonlinear relationships, tolerate correlated variables, provide high prediction accuracy, and objectively rank the importance of environmental factors. Implemented with RandomForestRegressor from sklearn [46].
@@
A gradient boosting framework optimized for efficiency and scalability. It builds decision trees sequentially, each one correcting the errors of the previous ones, using histogram-based methods and leaf-wise tree growth to speed up training [47]. In the context of Chl-a prediction with satellite imagery, LightGBM is particularly appealing because it can handle large volumes of pixel-level data efficiently, exploit subtle nonlinear interactions among spectral bands, and scale to the many candidate band combinations tested in our workflow. Implemented with LGMBRegressor from Microsoft’s LightGBM.
@@
A high-performance gradient boosting implementation that uses second-order derivatives for optimization, and includes regularization to avoid overfitting. Known for speed, accuracy and flexibility in hyperparameter tuning. For Chl-a prediction in optically complex waters, XGBoost is a strong candidate since it can model nonlinear relationships between reflectance and Chl-a, manage high-dimensional feature spaces (e.g. many band combinations), and incorporate regularization to maintain generalization across depths. Implemented with XGBRegressor from [48].
@@
A gradient boosting algorithm that is particularly designed to handle categorical features effectively, using techniques like ordered boosting and target statistics to reduce overfitting. CAT is included as a complementary approach to assess whether its regularization mechanisms and ordered boosting scheme improve generalization. Implemented with CatBoostRegressor [49].
@@
A feed-forward artificial neural network with one or more hidden layers. Each layer consists of neurons applying linear transformations followed by non-linear activation functions. The network is trained via backpropagation to minimize a loss function [50]. MLP represents the DL component of the comparison, capable of approximating highly nonlinear reflectance–chlorophyll relationships that may not be captured by tree-based or linear models. Implemented with MLPRegressor from sklearn.
@@
A meta-model that combines predictions from multiple base models (e.g. averaging, stacking) in order to improve predictive performance and robustness, by leveraging their complementary strengths and reducing individual model’s error variance. In this instance, a Ridge Regression, implemented with sklearn, is used to build the ensemble.
III-C Buoy data
There is a set of 12 buoys located in the Mar Menor, shown in Figure 1 and located in the coordinates from Table III according to [51] . These buoys measure several water quality parameters, such as oxygen, salinity, Chl-a, temperature or turbidity.
Figure 1: Buoy distribution in the Mar Menor. Top left map of Spain extracted from [52]
TABLE III: Control Points with Latitude and Longitude

### v2

@@ 2.2 Related Work
where $R(\lambda_{i})$ corresponds to the different band reflectances. During the study period from May 2015 to May 2017, the best results were achieved using the index normalized with the green and blue bands, reaching an R² of 0.88. Further work from the same authors [erenaMonitoringCoastalLagoon2019] focused on obtaining maps using the empirical algorithm previously proposed, and they also studied the relationship between Chl-a and turbidity. The study revealed the intricate variability between these parameters, as they were not consistently correlated due to factors such as atmospheric conditions, water column movements, and phytoplankton distribution.
@@
In 2020, [jimeno-saezUsingMachineLearningAlgorithms2020] used ML models to predict the concentration of Chl-a in the Mar Menor lagoon based on physical and chemical water data. Two approaches were evaluated: multilayer neural networks and support vector regression (SVR). Parameters such as water temperature, pH, suspended solids, and turbidity were used as predictor variables. However, this study does not use remote sensing or spectral bands; it is based exclusively on in situ measured parameters, so there are no spectral formulas or satellite image use. Yet, the study emphasizes that ML models are especially useful for monitoring this type of environment.
@@
The literature regarding remote sensing with Sentinel 2 for Mar Menor starts to proliferate in 2021. [gomezNewApproachMonitor2021] proposed a methodology using ML and deep learning (DL) models to estimate surface (less than 0.5 m) Chl-a. The inputs included band combinations such as the normalized difference chlorophyll index (NDCI), the normalized difference turbidity index (NDTI), and $B3/B8$ ratios, among others. Images are resampled up to 60 meters, and the most important features in the optimal model are the $B3/B8$ ratio, NDTI, and bands 9 and 2. Additionally, feature selection was explored in three scenarios: all variables, removal of features with correlations greater than 0.75, and Principal Component Analysis (PCA). Overall, maintaining all variables yielded the best results.
@@
[caballeroUseSentinel2Landsat82022] combined the two previously mentioned satellites: Sentinel 2 and Landsat 8. Their study focused on a brief period of time in 2021 and they implemented atmospheric corrections with ACOLITE, a toolbox designed to correct Level 1 products over marine, inland and coastal waters. Then, they estimated Chl-a using the bio-optical algorithm OC3 [MODISChlorophyllNOAA] based on water inherent properties and applied the model to the whole lagoon. In the case of [lopez-andreuMonitoringSystemMar2022], rather than applying the model to each pixel to map the entire water surface, the best performing model was applied only to points where in situ measurements were taken. Then, an interpolation method such as distance weighting or kriging was used to map the rest of the water surface.
@@
In [yeleyMarMenorLagoon2022], SNAP was used to correct Sentinel 2 images with the set C2X of C2RCC. With the atmospherically corrected images, reflectances are grouped within a 3 x 3 pixel window. Then, Chl-a is estimated using empirical models in which different band combinations are tested to fit the line $y=ax+b$. The band combination that yielded the best results is:
@@
While previous studies focused only on surface chlorophyll (less than one meter), [gimenezEnhancingShallowWater2024] explored prediction across the entire water column. They used several ML and DL models along with Sentinel-3 imagery to predict Chl-a concentrations at various depths for each in situ measurement point. They used 21 OLCI spectral bands as input and feature selection based on statistical significance. The best results are achieved on the surface, and as the depth increases, metrics deteriorate.
The most recent study in the Mar Menor region [gomez-jakobsenMonitoringChlorophyllConcentration2025] focused on using MODIS, VIIRS, and Sentinel-3 sensors to monitor Chl-a. The study developed an empirical algorithm named BELA based on the red-to-green ratio, which is expressed by the following formula: $Chl=0.353RG^{3}+2.132RG^{2}+3.905RG+2.110$ where $RG=\log_{10}(R_{670}/R_{550})$. However, when we applied this relationship to Sentinel 2 reflectances and our data, the results were not satisfactory. This may be due to the significant differences between the sensors in this article and the MSI, as well as the atmospheric corrections we used.
Remote sensing has been also widely used in other areas with turbid and productive water bodies. [mishraNormalizedDifferenceChlorophyll2012] introduced the NDCI to estimate Chl-a in Mobile Bay and the Mississippi Delta with MERIS imagery adjusting the model with simulated data to cover a wide range of chlorophyll values. Building on this, [pahlevanSeamlessRetrievalsChlorophylla2020] also employed NDCI in combination with probabilistic-based networks, arguing that chlorophyll is only one of several components influencing reflectance. [wattelezStatisticalAlgorithmEstimating2016] uses MODIS and proposes the idea of using a polynomial with logarithms for low values of Chl-a and a SVM for Chl-a greater than 3 mg/m3 while [linsAssessmentChlorophyllaRemote2017] explores the use several satellites in a tropical estuarine lagoon in Brasil. Finally, focusing on the Ebro Delta in Spain, [fernandez-tejedorAccurateEstimationChlorophylla2022] applied the C2RCC and C2X processors for atmospheric correction of Sentinel 2 imagery and tested several band combinations using linear, simple polynomial, and logarithmic models for chlorophyll estimation.
Table 2 summarizes all the articles discussed in a compressed format, so the reader can easily compared the methods used, band combinations, chlorophyll ranges, etc.
Table 2: Related Work summary table. Units for RMSE and MAE are mg/m3
@@ 10 B1 to B12
RF, LBM, XGB -
Average
of 0 to 5 m
- No
[yeleyMarMenorLagoon2022] Mar Menor Sentinel 2 10
@@
[gimenezEnhancingShallowWater2024] Mar Menor Sentinel 3 300 21 OLCI Bands
@@
[fernandez-tejedorAccurateEstimationChlorophylla2022] Ebro Delta Sentinel 2 10
@@
While numerous studies have demonstrated the potential of remote sensing for monitoring Chl-a in turbid and optically complex waters, certain aspects remain underexplored, particularly in regions such as the Mar Menor. Open-source methodologies that integrate long-term datasets, cover a broad range of Chl-a concentrations, explore predictions across the whole water column, and include systematic comparisons of multiple modeling strategies are still relatively scarce. Moreover, although neural network-based tools such as C2RCC and C2X are increasingly used, their integration into customized regional workflows and their comparison with other modeling approaches are areas that merit further development.
In this study, we address these gaps by developing and openly releasing a methodology tailored for the retrieval of Chl-a in optically complex coastal waters, using nearly a decade of Sentinel 2 data at 10 m resolution. This approach systematically evaluates a wide range of band combinations and processing strategies. It integrates neural-network-based atmospheric corrections (C2RCC, C2X, and C2X-Complex) with an advanced suite of machine learning regressors, including CatBoost, XGBoost, SVR, and other ML algorithms, to optimize chlorophyll-a (Chl-a) estimation across the water column. This comprehensive analysis is conducted within a robust multi-objective optimization and cross-validation framework. Additionally, we provide clear guidance and visual examples–in Section 5 and in the Github repository [^2] https://github.com/Antonio-MI/mar-menor-chl–of how the best-performing models can be operationally applied to generate Chl-a maps, contributing to improved transparency, reproducibility, and practical utility in regional water quality monitoring.
3 Materials and methods
@@
3.1 Study area
The Mar Menor is a coastal lagoon separated from the Mediterranean Sea by La Manga, with an approximate area of 135 km2 and shallow waters, less than 7 meters deep. The region’s climate is classified as a dry Mediterranean climate, with an average annual temperature ranging from 17°C to 18°C. The summers are hot, with temperatures over 40ºC, and the winters are mild, with extremes in temperature mitigated by the influence of the nearby sea. The region experiences low rainfall, typically ranging from 270 millimeters to 350 millimeters per year [MarcoFisico]. The temperature of the water body is subject to variation according to the season. During the winter months, the temperature is approximately 15°C. In the spring, the temperature typically ranges from 16–18°C. During the summer months, the temperature reaches an average of 24-25°C, and there have been recorded temperatures of 31-32°C during recent episodes.
The salinity of the Mar Menor typically varies between 40 and 47 Practical Salinity Unit (PSU), which is considered high for a coastal lagoon and attributed to minimal rainfall and substantial evaporation, and classifies it as an hypersaline coastal lagoon. In hydrodynamic terms, the Mar Menor displays marked homogeneity in its internal behavior, with minor variations in level between sensors except during episodes of intense wind, where seiche-type oscillations are generated in the lagoon. The wind also generates a water level difference of up to 10 centimeters and directs surface currents in its direction, thereby facilitating water mixing and renewal [InformeIntegralSobre2017].
3.2 Methods
This subsection outlines the ML models that were implemented. This selection explores a wide range of algorithms, covering classical statistical approaches and modern ML techniques. Some of these models were selected because previous studies have already demonstrated their suitability for predicting Chl-a and related water quality parameters in coastal and inland waters. Others were included to expand upon those findings and assess whether additional modeling paradigms could improve robustness and generalization. This includes distance-based methods, linear and regularized regression models, tree-based learners, and gradient boosting frameworks. Additionally, a simple multilayer perceptron is also included, and ensemble strategies are used to combine individual predictors. This variety of models enables a thorough comparison of different learning paradigms and demonstrates their respective strengths in Chl-a prediction.
@@ 3.2 Methods
A supervised learning method in which the output is modeled as a linear combination of input features, capturing the approximately linear relationships between Chl-a concentration and spectral reflectance [matus2018predictive]. The model assumes a constant rate of change in the output for each unit change in the inputs, and its parameters are typically estimated by minimizing the sum of squared errors [LinearRegressionML]. Implemented with LinearRegression from sklearn.
@@
A non-parametric regression method where the prediction for a given instance is based on the average of the target values of its $k$ closest neighbors in the feature space. In this work the weights are assigned proportionally to the inverse distance from the query point, and distance is computed as euclidean distance. KNN could suit this problem because it can capture local nonlinearities between reflectance and Chl-a concentration without assuming a specific functional form. Implemented with KNeighborsRegressor from sklearn [KNeighborsRegressor].
@@
A regression algorithm derived from Support Vector Machines (SVM) that seeks to fit a function within an error margin (an epsilon-insensitive tube) while penalizing points that fall outside this margin. SVR effectively captures complex nonlinear relationships between environmental factors and Chl-a concentration, avoids overfitting through structural risk minimization, performs well with small sample sizes, and provides high prediction accuracy even when data are noisy [xu2023prediction]. Implemented using the SVR module from scikit-learn [SVR].
@@
A linear regression model that combines L1 (lasso) and L2 (ridge) regularization penalties. It is useful when there are many correlated features: L1 helps with feature selection (sparse coefficients), L2 stabilizes the coefficient values. In this context, ELN helps reduce redundancy between highly correlated spectral bands while maintaining model stability. Implemented with ElasticNet from sklearn [ElasticNet].
Random Forest (RF) An ensemble learning method that builds many decision trees on bootstrapped subsets of the data and averages their predictions to reduce overfitting and improve generalization. RF can handle nonlinear relationships, tolerate correlated variables, provide high prediction accuracy, and objectively rank the importance of environmental factors. Implemented with RandomForestRegressor from sklearn [RandomForestRegressor].
@@
A gradient boosting framework optimized for efficiency and scalability. It builds decision trees sequentially, each one correcting the errors of the previous ones, using histogram-based methods and leaf-wise tree growth to speed up training [LightGBM]. In the context of Chl-a prediction with satellite imagery, LightGBM is particularly appealing because it can handle large volumes of pixel-level data efficiently, exploit subtle nonlinear interactions among spectral bands, and scale to the many candidate band combinations tested in our workflow. Implemented with LGMBRegressor from Microsoft’s LightGBM.
@@
A high-performance gradient boosting implementation that uses second-order derivatives for optimization, and includes regularization to avoid overfitting. Known for speed, accuracy and flexibility in hyperparameter tuning. For Chl-a prediction in optically complex waters, XGBoost is a strong candidate since it can model nonlinear relationships between reflectance and Chl-a, manage high-dimensional feature spaces (e.g. many band combinations), and incorporate regularization to maintain generalization across depths. Implemented with XGBRegressor from [XGBoost].
@@
A gradient boosting algorithm that is particularly designed to handle categorical features effectively, using techniques like ordered boosting and target statistics to reduce overfitting. CAT is included as a complementary approach to assess whether its regularization mechanisms and ordered boosting scheme improve generalization. Implemented with CatBoostRegressor [CatBoost].
@@
A feed-forward artificial neural network with one or more hidden layers. Each layer consists of neurons applying linear transformations followed by non-linear activation functions. The network is trained via backpropagation to minimize a loss function [MLPRegressor]. The MLP is included with the intention of approximating highly nonlinear reflectance–chlorophyll relationships that may not be captured by tree-based or linear models. Implemented with MLPRegressor from sklearn.
@@
A meta-model that combines predictions from multiple base models (e.g. averaging, stacking) in order to improve predictive performance and robustness, by leveraging their complementary strengths and reducing individual model’s error variance. In this instance, the ensemble is built by averaging the other models.
3.3 Buoy data
There is a set of 12 buoys located in the Mar Menor, shown in Figure 1 and located in the coordinates from Table 3 according to [tomadedatosenelmar] . These buoys measure several water quality parameters, such as oxygen, salinity, Chl-a, temperature or turbidity.
Figure 1: Buoy distribution in the Mar Menor. Top left map of Iberian Peninsula extracted from [ORSMaps]
Table 3: Control Points with Latitude and Longitude